# Enhancing End Stage Renal Disease Outcome Prediction: A Multi-Sourced Data-Driven Approach

### v3

@@ Abstract
Objective: To improve prediction of Chronic Kidney Disease (CKD) progression to End Stage Renal Disease (ESRD) using machine learning (ML) and deep learning (DL) models applied to an integrated clinical and claims dataset of varying observation windows, supported by explainable AI (XAI) to enhance interpretability and reduce bias.
Materials and Methods: We utilized data from 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018. Following data preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using data extracted from five distinct observation windows. Feature importance and SHAP analysis were employed to understand key predictors. Models were tested for robustness, clinical relevance, misclassification errors and bias issues.
Results: Integrated data models outperformed those using single data sources, with the Long Short-Term Memory (LSTM) model achieving the highest AUC (0.93) and F1 score (0.65). A 24-month observation window was identified as optimal for balancing early detection and prediction accuracy. The 2021 eGFR equation improved prediction accuracy and reduced racial bias, especially for African American patients.
Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities for this population.
Conclusion: This study presents a robust framework for predicting ESRD outcomes in CKD patients, improving clinical decision-making and patient care through multi-sourced, integrated data and AI/ML methods. Future research will expand data integration and explore the application of this framework to other chronic diseases.
@@ 1 Introduction
Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 5 or administrative claims data 6; 7. While informative, these studies often utilize a limited set of features, potentially under-representing the complexity of CKD progression. For instance, Sharma et al. 8 developed a model using claims data to identify CKD patients at risk of Hyperkalemia, and Krishnamurthy et al. 6 predicted CKD onset using comorbidities and medications from Taiwan’s National Health Insurance Database. However, claims-based studies generally lack the granularity of clinical data, which offers detailed insights through laboratory results, medication records, and patient demographics. For example, Tangri et al. 9 used clinical variables such as age, gender, estimated glomerular filtration rate (eGFR)10, and albuminuria to predict ESRD progression, and Keane et al. 11 utilized serum creatinine and urine protein levels for high-risk patient identification. Despite their value, models based solely on clinical data may miss the full spectrum of patient interactions with the healthcare system and often struggle with missing data and inconsistent recording practices. Additionally, they may overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse patient populations.
In the field of explainable AI (XAI) for healthcare, established methods such as SHapley Additive exPlanations (SHAP) 12 and Local Interpretable Model-agnostic Explanations (LIME) 13 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models. However, relying solely on a single source of data — whether claims or clinical — restricts the ability to fully understand the complexity and variability of CKD progression across a diverse patient population and predict ESRD, thereby limiting the real-world applicability of these models.
This study aims to bridge this gap by developing a robust and comprehensive framework that not only utilizes clinical or claims data in isolation but also harnesses the combined strengths of these integrated data sets. At the same time, by minimizing the observation window needed for accurate prediction, our framework integrates clinical relevance with a practical, patient-centered focus. This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making to ultimately improve patient outcomes.
@@ 2 Objective
This research aims to evaluate prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data. Fig. 1 illustrates the key temporal components of our study design, with the observation window beginning at Stage 3 diagnosis (t=0). During this observation window, patients may progress through different CKD stages (S4, S5); however, we specifically focus on the cohort that does not develop ESRD within the observation window itself. This ensures that our predictive modeling uses only pre-ESRD data to forecast future ESRD development. Our primary objectives are:
  • ESRD Risk Prediction: To estimate the probability that a CKD patient will develop ESRD in the future, based on data from a specified observation window that begins from each patient’s first occurrence of CKD stage 3. This probability is represented as:
@@
  • Observation Window Optimization: To identify the optimal observation window $T_{\mathrm{obs}}$ that provides the best performance across all candidate windows $t_{\mathrm{obs}}$ (6, 12, 18, 24, and 30 months), ensuring accurate predictions while minimizing the length of the observation window required:
@@
    where Performance $(t)$ is evaluated based on key predictive metrics: F1-score (the harmonic mean of precision and recall) 14, Area Under the Receiver Operating Characteristic curve (AUROC) 15, and Area Under the Precision-Recall Curve (AUPRC) 16.
Figure 1: A timeline illustration of observation windows for CKD progression to ESRD. The observation starts at CKD Stage 3. Objective 1 estimates the probability of ESRD occurrence using clinical and claims data within the observation window, while Objective 2 identifies the optimal observation window that maximizes predictive performance.
This approach balances the trade-off between maximizing predictive accuracy and minimizing the duration of the observation window required for reliable predictions. Additionally, we apply explainable AI techniques, such as feature importance and SHAP analysis, to enhance model transparency at both the cohort level to identify key factors, and the individual level to provide personalized insights.
@@ 3 Data Sources
This study utilized two key de-identified datasets from different healthcare organizations: administrative claims data and clinical data. The de-identified administrative claims dataset was provided by Highmark, a major health insurance provider, and includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018. This dataset encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with CKD. The use of this dataset was approved by Highmark’s Institutional Review Board.
The clinical dataset was sourced from the EHR data of TMA, P.C., a leading community nephrology practice. This dataset contains laboratory results, patient demographics, diagnostic details, and medication records spanning from 1998 to 2018. For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data (2009–2018). Usage of the clinical data was approved by the Institutional Review Board of Carnegie Mellon University.
Preprocessing of the claims data involved removing duplicates, excluding records without any CKD stage information, and eliminating entries with negative costs to ensure data integrity. After cleaning, the claims dataset included 5,317,178 individual claims across 7,129 unique patients. The clinical dataset was similarly refined to include 433,421 laboratory records for 10,326 patients. These datasets were then integrated using unique patient IDs to provide a comprehensive view of each patient’s medical and claims history.
@@ 4 Methods
As shown in Figure 7, our approach to predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) encompasses three primary stages: Data Preparation, Modeling and Validation, and Additional Analyses to assess robustness and clinical utility.
Figure 2: Overview of research framework and training strategy. Part A provides an overview of the data preparation, modeling and validation, and additional analyses pipeline. Part B illustrates the cohort identification process, detailing the inclusion and exclusion criteria applied to claims and clinical records. Part C outlines the modeling strategy.
@@ 4.1.1 Data Pre-processing
As mentioned earlier, we performed data cleansing to ensure the integrity of the datasets. This process involved removing duplicate records to prevent redundancy and potential biases in the analysis. We also excluded entries lacking any stage of CKD diagnoses to maintain the datasets’ relevance to our study objectives, and claims records with negative values, which likely indicated data entry errors.
To address missing data issue, we employed Multiple imputation using chained equation (MICE)17, a widely adopted method in healthcare research for its ability to account for multivariable relationships and provide robust imputations. Normalization was applied to ensure that features contributed proportionately to the analysis. We scaled numerical values to a standard range, a step essential for algorithms sensitive to the magnitude of input values. We applied log transformations 18 to variables exhibiting skewed distributions to stabilize variance and approximate normality. This approach is particularly beneficial for handling data with exponential growth patterns, as it mitigates the impact of outliers and enhances the performance of linear models.
In the laboratory data, we observed an additional challenge where many records lacked unit specifications, leading to significant discrepancies in numerical values. To address this, we conducted distribution detection & adjustment19 to identify and correct unit inconsistencies. By analyzing the statistical distribution of each laboratory measurement, we detected shifts indicative of differing units. In consultation with clinical experts, we established expected value ranges for each measurement. Outlier values that deviated markedly from these ranges were adjusted to align with the appropriate units, ensuring consistency and comparability across all records.
@@ 4.1.2 Cohort Identification
In Part B of Fig.7, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period. Although the example in the figure uses an 18-month observation window, the exact duration may vary based on the selected window of observed data.
The study cohort was further curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists. We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management, resulting in a final cohort with uninterrupted, detailed clinical and claims histories.
@@ 4.1.3 Feature Engineering
Part C depicted in Fig. 7 involves identifying and creating predictive variables essential for modeling. With claims data, we categorized the features used for prediction into two main groups. The first, cost-based features, was derived from five types of claims: Inpatient, Outpatient, Professional, Pharmacy, and Vision. These features include the count of unique claims per type, the aggregate cost of each type, the range of claim costs, and the standard deviation of claim costs for each patient.
The second group, comorbidity-based features, includes patient-specific clinical information such as the patient duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia.
Clinical data features were categorized into three groups based on consultations with nephrologists, commonly referenced literature, and the characteristics of our datasets. These groups include demographic features, laboratory data features, and comorbidity features.
  • Demographic Features: These include age, gender, ethnicity and body mass index (BMI), which have been widely used for understanding CKD progression.
  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, and bicarbonate, which are important for monitoring CKD. Notably, urine albumin-to-creatinine ratio (UACR) 20 and sodium were not selected due to insufficient readings in the dataset.
  • Additional Comorbidity Features: These capture the presence of additional health conditions such as cardiovascular disease, anemia, and metabolic acidosis, in addition to those identified using claims data, and have been shown to influence CKD outcomes.
Following these steps, statistical analyses were performed to compare variables between patients who progressed to ESRD and those who did not. For continuous variables, data were assessed for normality; those meeting normality assumptions were analyzed using independent t-tests to compare group means. For non-normally distributed variables, logarithmic transformations were applied during preprocessing to approximate normality before conducting t-tests. Categorical variables were expressed as frequencies (percentages) and compared using chi-squared tests. Statistical significance was set at $p<0.05$ for all analyses.
@@ 4.2 Modeling & Validation
To develop predictive models for CKD progression to ESRD, we employed both machine learning (ML) and deep learning (DL) frameworks to capture complex data patterns. Our overarching strategy involved partitioning the dataset into training, validation, and testing subsets through stratified sampling, maintaining class distribution across all splits. This approach is particularly crucial in medical research, where class imbalances can significantly impact model performance. To further address the imbalance, we applied SMOTE 21; 22 to the training data, generating synthetic minority samples and enhancing the model’s ability to learn from underrepresented classes.
@@ 4.2.1 Machine Learning Methods
We employed statistical methods, beginning with logistic regression (LR) 23 as a baseline model to estimate the probability of ESRD progression using clinical and claims data features. Beyond LR, we utilized ML techniques such as Random Forest (RF) 24 and Extreme Gradient Boosting (XGBoost) 25. We evaluated performance using a standard k-fold cross-validation setup 26. RF uses ensemble learning by constructing multiple decision trees, which improves model accuracy and mitigates overfitting. XGBoost, a gradient boosting algorithm, builds models iteratively to correct errors from previous iterations, proving highly effective for structured data.
To enhance model interpretability, we conducted cohort-level feature importance analysis to identify key predictors across the dataset. Additionally, SHAP analysis was applied at the individual level to elucidate each feature’s contribution to specific predictions. This dual approach—cohort-level and individual-level analysis—not only provides insights into the drivers of ESRD progression but also improves model transparency, supporting personalized treatment decisions.
@@ 4.2.2 Deep Learning Methods
To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 7; 27 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis. These three-month intervals are indexed as timestamps (timestamp 0 for months 0-3, timestamp 1 for months 3-6, and so on), with a 18-month observation window containing six timestamps in total. This temporal segmentation approach serves multiple purposes: it captures the temporal evolution of CKD, reflects its gradual progression, and aligns with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data. Numerical features were aggregated, categorical features were encoded binarily to capture conditions or events, and a baseline assessment of historical records was conducted to adjust for pre-existing conditions, ensuring the model accurately accounted for CKD presence and progression. This timestamp-based structure enables our model to track the progression of disease markers and identify critical temporal patterns that may indicate increased risk of ESRD development.
We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 28 for feature extraction, Recurrent Neural Network (RNN) 29, Long Short-Term Memory network (LSTMs 30, Gated Recurrent Unit (GRU) 31 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 32 for sequence modeling. These models were chosen to provide a robust foundation for accurately predicting ESRD progression, identifying key risk factors, and anticipating CKD to ESRD transition with high precision. We forgo k-fold cross-validation due to high computational costs and training variance from random initialization33. Modern regularization techniques like dropout reduce this need, so we opted for stratified holdout validation to balance efficiency and reliability while maintaining class distributions.
4.2.3 Hyperparameter Tuning
For performance optimization, we employed specialized Python libraries (Python 3.10) to optimize model performance through hyperparameter tuning. For machine learning models, we utilized the H2O framework 34, which provides comprehensive grid search capabilities. The framework systematically explores a predefined set of hyperparameters to identify the optimal combination for model performance. For deep learning models, we leveraged Microsoft’s Neural Network Intelligence (NNI) 35. NNI is an open-source Python toolkit that automates neural architecture search and hyperparameter optimization. It efficiently dispatches and manages trial jobs generated by various tuning algorithms, facilitating the search for optimal neural architectures and hyperparameter configurations.

### v4

@@ Abstract
Objective: To improve prediction of Chronic Kidney Disease progression to End Stage Renal Disease using machine learning and deep learning models applied to integrated clinical and claims data with varying observation windows, supported by explainable AI to enhance interpretability and reduce bias.
Materials and Methods: We utilized data from 10,326 CKD patients, combining clinical and claims information from 2009-2018. After preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using five distinct observation windows. Feature importance and SHAP analysis were employed to understand key predictors. Models were tested for robustness, clinical relevance, misclassification patterns, and bias.
Results: Integrated data models outperformed single data source models, with Long Short-Term Memory (LSTM) achieving the highest AUC (0.93) and F1 score (0.65). A 24-month observation window optimally balanced early detection and prediction accuracy. The 2021 eGFR equation improved prediction accuracy and reduced racial bias, particularly for African American patients.
Discussion: Improved prediction accuracy, interpretability and bias mitigation strategies have the potential to enhance CKD management, support targeted interventions, and reduce healthcare disparities.
Conclusion: This study presents a robust framework for predicting ESRD outcomes, improving clinical decision-making through integrated multi-sourced data and advanced analytics. Future research will expand data integration and extend this framework to other chronic diseases.
@@ 1 Introduction
Previous CKD progression prediction efforts used either EHR clinical data 5 or administrative claims 6; 7. These approaches often use limited features, not fully capturing CKD progression complexity. Sharma et al. 8 developed a claims-based model to identify CKD patients at Hyperkalemia risk, while Krishnamurthy et al. 6 predicted CKD onset using comorbidities and medications. Claims data typically lacks clinical data’s granularity. Tangri et al. 9 used age, gender, and eGFR 10 to predict ESRD progression, and Keane et al. 11 utilized serum creatinine and urine protein levels for high-risk patient identification. Models using only clinical data may miss complete healthcare system interactions, struggle with missing data and inconsistent recording, and overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse populations.
This study bridges a critical gap by developing a framework that utilizes integrated clinical and claims data rather than isolated data sources. By minimizing the observation window needed for accurate predictions, our approach balances clinical relevance with patient-centered practicality. This integration enhances both predictive accuracy and clinical utility, enabling more informed decision-making to improve patient outcomes.
@@ 2 Objective
This research evaluates predictive models for CKD progression to ESRD using integrated clinical and claims data. Fig. 1 illustrates our study design, where the observation window begins at the initial diagnosis of CKD Stage 3 (t=0). Although patients may progress through subsequent CKD stages (Stages 4 and 5) during this observation period, our cohort specifically excludes those developing ESRD within this timeframe. This ensures that predictive modeling utilizes only pre-ESRD data for forecasting future ESRD onset.
  • Our primary objective, ESRD risk prediction, aims to estimate the probability that a patient diagnosed with CKD Stage 3 will develop ESRD after the observation window, represented as follows:
@@
  • To identify the optimal observation window $T_{\mathrm{obs}}$ that provides the best performance across all candidate windows $t_{\mathrm{obs}}$ (6, 12, 18, 24, and 30 months), ensuring accurate predictions while minimizing the length of the observation window required:
@@
    where Performance $(t)$ is evaluated based on key predictive metrics: F1-score 12, Area Under the Receiver Operating Characteristic curve (AUROC) 13, and Area Under the Precision-Recall Curve (AUPRC) 14.
Figure 1: A timeline illustration of observation windows for CKD progression to ESRD. The observation starts at CKD Stage 3. Objective 1 estimates the probability of ESRD occurrence after the observation window, using clinical and claims data within the observation window, while Objective 2 identifies the optimal observation window that maximizes predictive performance.
@@ 3 Data Sources
This study utilized two integrated datasets: administrative claims data and clinical data. The claims dataset includes patient healthcare interactions, spanning from 2009 to 2018, containing diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with CKD. The clinical dataset from electronic health records contains laboratory results, patient demographics, diagnostic details, and medication records, which was truncated to match the 10-year span of the claims data for consistency of integration. See Appendix A for data details.
@@ 4 Methods
As illustrated in Figure 2(a), our comprehensive methodology for predicting CKD progression to ESRD consists of three primary stages: Data Preparation, Modeling, and Additional Analyses. Within the Modeling stage, we employ a two-phase analytical framework that first systematically evaluates various predictive models across multiple observation windows (but the same patient cohort for consistent comparisons) to determine optimal performance, and subsequently retrains the most effective model on the cohort for the each observation window for practical clinical application. This framework bridges statistical rigor and clinical relevance, facilitating actionable insights for nephrologists managing patients with CKD stage 3 and above.
(a) Three-stage pipeline for predicting CKD progression to ESRD. Stage 1: feature engineering, data preprocessing, and cohort identification using clinical and claims data. Stage 2: (2A) systematic evaluation of observation windows with ML/DL models; (2B) retraining the optimal model on the filtered cohort from the optimal observation window, followed by explainable AI analysis. Stage 3: Additional analyses including bias detection and misclassification assessment.
(b) Multi-step cohort curation process for longitudinal CKD analysis. Starting with initial selection of patients at CKD stage 3, an adjustable observation window (18 months shown) was applied. Only patients with complete clinical and claims data throughout this period were retained, while those progressing to ESRD within the observation window were excluded, ensuring a cohort for analysis (highlighted in yellow) with uninterrupted longitudinal records.
Figure 2: Framework for CKD progression prediction: (A) Overview of the three-stage analytical pipeline, and (B) Cohort identification methodology.
@@ 4.1.1 Data Pre-processing
We removed duplicate records to prevent redundancy and bias in the analysis. Entries lacking CKD diagnoses were excluded to maintain relevance to study objectives, along with claims containing negative values (likely data entry errors). The cleaned claims dataset comprised 5,317,178 claims across 7,129 unique patients, while the refined clinical dataset included 433,421 laboratory records for 10,326 patients. These datasets were integrated using unique patient IDs to provide a comprehensive view of each patient’s medical and claims history.
To address missing data, we employed multiple imputation via chained equations (MICE) 15. Continuous numerical features were standardized, and skewed variables were log-transformed 16 to stabilize variance and mitigate outliers. For laboratory values missing unit specifications, we used distribution detection and adjustment 17 to identify outliers and harmonize units across records (medical expert-advised ranges in Appendix B).
@@ 4.1.2 Cohort Identification
As depicted in Fig. 2(b), cohorts were curated through a multi‐step process: patients at CKD stage 3 were selected and an observation window—18 months in the example, but adjustable per study—was applied. Only individuals with complete clinical and claims data throughout this window were retained, and those who progressed to ESRD within it were excluded, yielding a cohort for analysis.
@@ 4.1.3 Feature Engineering
Feature engineering involved identifying predictive variables essential for modeling. For claims data, we derived two main groups of features: cost-based features, including claim counts, aggregate patient costs, cost ranges, and cost standard deviations from inpatient, outpatient, professional, pharmacy, and vision claims; and comorbidity-based features, such as CKD stage 3 duration, emergency department visit frequency, and critical comorbidities (e.g., hypertension, diabetes, phosphatemia).
Clinical data features were categorized based on nephrology expert consultations, relevant literature, and dataset characteristics into demographic features, including age, gender, ethnicity, and BMI as established CKD progression indicators; laboratory features comprising essential clinical parameters like eGFR, hemoglobin, phosphorus, serum calcium, and bicarbonate (excluding UACR 18 and sodium due to insufficient data availability); and additional comorbidity features, capturing conditions such as cardiovascular disease, anemia, and metabolic acidosis, which are known to influence CKD outcomes.
Statistical analyses compared variables between patients progressing to ESRD and those who did not. For continuous variables, data were assessed for normality; normally distributed variables were analyzed using independent t-tests to compare group means. Non-normally distributed variables underwent logarithmic transformation during preprocessing before applying t-test. Categorical variables were expressed as frequencies (percentages) and compared using chi-squared tests. Statistical significance was set at $p<0.05$ for all analyses.
@@ 4.2 Modeling & Validation
We trained both ML and DL models on stratified train/validation/test splits to preserve class proportions, and applied SMOTE 19; 20 to the training data to oversample the minority class.
@@ 4.2.1 Machine Learning Methods
We employed logistic regression (LR) 21 as our baseline model for ESRD progression probability estimation. We extended our analysis to Random Forest (RF) 22 and XGBoost 23, evaluating performance via k-fold cross-validation 24. RF utilizes multiple decision trees to enhance accuracy while reducing overfitting. XGBoost sequentially builds models to correct previous errors, proving highly effective for structured data.
@@ 4.2.2 Deep Learning Methods
To develop a robust predictive model for CKD progression to ESRD, we adopted methodologies from prior studies 7; 25 by constructing a longitudinal data representation with three-month intervals following initial CKD stage 3 diagnosis. These intervals were sequentially indexed as timestamps (0 for months 0-3, 1 for months 3-6, etc.).
This temporal segmentation approach serves several purposes, capturing CKD’s temporal evolution, reflecting its gradual progression, aligning with clinical guidelines recommending periodic nephrologist visits, and accommodating potential time lags between clinical events and their corresponding claims data. These lags, particularly notable for critical events such as transplantation or dialysis initiation, exhibit substantial variability both across and within patient records, complicating consistent correction strategies. Employing three-month intervals mitigates the impact of these temporal discrepancies, improving model accuracy. Numerical features were aggregated, categorical features encoded binarily to reflect conditions or events, and baseline historical records were analyzed to adjust for pre-existing conditions, thereby ensuring accurate representation of CKD progression. This timestamp-based structure allows models to capture disease progression dynamics and identify critical temporal patterns signaling heightened ESRD risk.
We evaluated multiple deep learning architectures, including Convolutional Neural Networks (CNN)26 for feature extraction; Recurrent Neural Networks (RNN)27, Long Short-Term Memory networks (LSTM)28, and Gated Recurrent Units (GRU)29 for modeling temporal dependencies; and Temporal Convolutional Networks (TCN)30 for sequence modeling. These architectures were selected to optimize prediction accuracy of ESRD progression and improve identification of associated risk factors. Detailed implementation information is available in Appendix D.
To maximize clinical relevance, our analytical framework employs a two-phase approach (as illustrated in Fig. 2(a) Stage 2). First, we systematically compare all modeling techniques mentioned above across multiple observation windows to identify optimal predictive performance. Second, recognizing that clinicians cannot predetermine a patient’s progression timeline to ESRD in real-world settings, we re-train the best-performing model using the complete cohort with the optimal observation window. This clinically-oriented model then undergoes comprehensive explainable AI analysis to identify key predictive features and temporal patterns. This approach bridges the gap between statistical performance and practical clinical implementation, providing actionable insights for nephrologists.
4.3 Explainable AI Methods
After identifying the best-performing model and optimal observation window, we employed two complementary explainable AI techniques to understand the features driving these predictions. At the cohort level, we utilized feature importance analysis to identify key predictive variables across the population. At the individual level, we applied SHAP analysis to provide detailed, patient-specific insights into model predictions.